# The Landscape of GPU-Centric Communication

### v2

@@ 1. Introduction
In recent years, GPUs have become the accelerator of choice for a vast array of applications across both HPC and ML. This quick adoption, driven mainly by the GPU’s massive parallelism and fast memory bandwidth, means that most of modern cloud and HPC capability is now concentrated in clusters of GPUs. As of June 2024, 9 of the 10 leading Top500 supercomputers rely on GPU clusters for acceleration and this trend is likely to continue (Meuer et al., 2023). The only one that does not rely on GPUs relies on highly-vectorized CPU architecture with High-Bandwidth Memory.
While using scores of GPUs has been shown to significantly accelerate computation, communication between the GPUs can quickly become a scalability bottleneck (Shimokawabe et al., 2011). Traditionally, multi-GPU communication, both within and across nodes, had always been the responsibility of the CPU. From their conception, GPUs were thought of as devices that can supply large amounts of computation but that are also inherently dependent on the CPU for auxiliary tasks like communication. In this CPU-centric model of execution, the routines relaying data for consumption by the GPUs are oblivious to the GPUs’ existence.
@@
  • In Section 2, we define the terminology, provide a definition for GPU-centric communication and taxonomize the existing approaches within a node and across nodes to eliminate any confusion.
  • In Section 3, we discuss the vendor-provided mechanisms for enabling communication and networking, managing memory across devices in a multi-GPU execution. These mechanisms are used as building blocks for higher-level GPU-centric software libraries.
  • In Section 4, we list the main communication libraries for both intra- and inter-node setups, discuss their benefits and challenges and rely on existing benchmarking papers to provide insights into their performance.
@@
We also acknowledge that certain methods and technologies depend on proprietary ecosystems. In this work, we primarily focus on NVIDIA solutions. However, similar solutions are offered by AMD, and we explicitly highlight those that are only offered by one of the two vendors. Intel also offers high-end HPC/ML GPUs. However, at the time of writing, there is not much public information available about the outlook of the product line. Intel Ponte Vecchio is not going to be deployed on new clusters (the Home, 2024), and the next generation Falcon Shores GPUs are not going to be released before late 2025 (Shah, 2024).
@@ 2. Terminology and Communication Types
Type API Data Path Examples
3 Device native Device Device NVSHMEM/ROCSHMEM GPU-side, direct load/store (P2P)
4 Host fallback Device Host Direct load/store (No P2P)
Table 1. Types of intra-node communication methods
We can loosely define GPU-centric communication as mechanisms that reduce the involvement of the CPU in the critical path of multi-GPU execution. This is a very broad definition covering a wide spectrum of solutions, involving both the vendor-level improvements that grant GPUs autonomy in communication and user-level implementations that leverage those improvements. To make this distinction clear, we discuss them in separate sections. In Section 3, we focus on communication mechanisms and primitives provided natively as part of the NVIDIA CUDA and AMD ROCm runtimes. In Section 4, we discuss how those mechanisms give rise to higher-level GPU-centric communication libraries.
@@ 4 Host fallback Device Host Direct load/store (No P2P)
The communication method in 1, referred as host native, is made on the host side and does not involve direct P2P (peer-to-peer) access between the devices. These involve all methods that can be launched on the host side with P2P access disabled. Otherwise, as 2 (host-controlled) shows, the communication does not involve an extra copy to the host memory, and passes directly through PCIe, NVLink or Infinity Fabric interconnects. NCCL, GPU-aware MPI and *memcpy operations have host-side API, so they may belong to both 1 and 2. With direct access to peer devices memory the device-side API allowed to remove the CPU both from the data path and control path in intra-node communication, as 3, referred as device native, shows on Figure 1. Both NVSHMEM and ROCSHMEM offer host-side API as well, but require P2P access, so its host-side API belongs to 2 and device-side API is of type 3. In-kernel P2P direct load and stores offer similar functionality and belong to type 3, but work even with P2P access disabled, which is type 4, where data path falls back to the host.
@@ 1 Host native Host Host Host Through host (2 copies) GPU-Aware MPI before 2
2 GPUDirect 1.0 Host Host Host Through host (1 copy) GPU-Aware MPI, NCCL
3 GPUDirect RDMA, ROCnRDMA D/H Host Host Direct GPU-Aware MPI, NCCL, NVSHMEM
4 GPUDirect Async D/H Host Device Depends on 3 GPU-Aware MPI, NVSHMEM
@@ 5 Device native Device Device Device Depends on 3 NVSHMEM with IBGDA, GPUrdma
We identify five main categories when classifying inter-node communication. From 1 to 5, as Table 2 and Figure 2 show, more components of communication calls were moved to the GPU side, while the data transfer path saw a reduction in the number of data copies required to reach the NIC. Over the years, the communication methods and the corresponding technologies depict the optimizations of both data transfer and communication control, which will be discussed in Section 3. First, entirely CPU-side communication method in 1 was available, which has been improved by removal of extra copy between CPU-GPU and CPU-NIC buffers in 2. After that, a GPU-side optimization facilitated direct access to the GPU memory from NIC 3 minimizing the data path between them. 4 represents GPU-triggered communication technologies such as GPUDirect Async and GPU-TN (LeBeane et al., 2017), where the GPU has become capable to initiate communications, given that CPU prepared the packets on the NIC in advance. 5 moves the packet preparation and interaction with the NIC entirely to the GPU as well, making device-native communication possible.
The types given in Table 2 and Figure 2 do not reflect all possible combinations, since some libraries, based on the configuration and the available hardware may lead to different combinations of data paths and control. For example, without the RDMA technology, even with GPU-side communication control, the data path will involve the host memory.
@@ 3. Vendor Mechanisms
In this section we discuss the vendor-provided mechanisms for enabling communication and networking, managing memory across devices in a multi-GPU execution. These mechanisms are provided by GPU programming model runtimes or as part of the extended APIs.
@@
Figure 3 summarizes the technologies provided by NVIDIA, detailing their timeline and availability. The technologies are classified into four categories: memory managers, GPUDirect technologies, hardware, and libraries. Next, we introduce the memory management mechanisms and GPUDirect technologies, followed by the hardware support that served as precursors and ultimately made these communication methods viable. These technologies form the backbone of the higher-level GPU-centric libraries, which are discussed in Section 4.
@@ 3.1.1. Page-Locked / Pinned Memory
By default, memory allocated on the host using cudaMalloc() is pageable and is not accessible to the GPU. When a transfer between pageable host memory and device memory is performed, the GPU runtime must first stage the host data through a temporary buffer in page-locked memory and then copy the data from page-locked memory to the GPU. To avoid the pageable $\rightarrow$ page-locked memory copy, cudaMallocHost() allows allocating page-locked memory directly, skipping the intermediate copy stage. Because of this, page-locked memory is also referred to as zero-copy or pinned memory (NVIDIA, 2023a; Harris, 2012).
Pinned memory, known for its high bandwidth and low latencies in host-device transfers (Pearson et al., 2019), efficiently coordinates CPU-GPU execution by enabling direct system-wide access. It is also utilized with GPUDirect RDMA for improved inter-node communication (Li et al., 2020). However, its physical memory locking can lead to significant memory consumption, potentially impacting system performance with excessive allocations (Harris, 2012).
@@ 3.1.2. Unified Virtual Addressing (UVA)
UVA is a memory management technique introduced in CUDA 4.0 which allows all GPUs and CPUs within a node to share the same unified virtual address space (NVIDIA, 2023a, 2011). Prior to UVA, host $\leftrightarrow$ device and device $\leftrightarrow$ device copies had to explicitly specify the direction of transfer. With UVA, the physical memory location can be inferred from pointer values, thus, reducing the overhead of managing separate memory spaces and enabling libraries to simplify their interfaces (Schroeder, 2011).
3.1.3. CUDA IPC
In early CUDA versions, pointers could not be accessed across process boundaries, so memory copies between GPU buffers had to go through the host, creating a bottleneck. To overcome this limitation, CUDA 4.1 introduced CUDA Inter-Process Communication (IPC), which enables processes on the same node to access device buffers of other processes without additional copies (NVIDIA, 2017a). With CUDA IPC, memory handles are created and passed between processes using standard IPC mechanisms, resulting in lower latencies than staging copies through the host. However, the overhead of creating memory handles can be significant and may offset the latency benefits (Potluri et al., 2012).
@@ 3.1.4. Unified Virtual Memory (UVM)
Introduced in CUDA 6.0, UVM allows for the allocation of managed memory through cudaMallocManaged() calls by creating a single address space accessible to all processors within a single node. UVM works by dividing the requested memory into pages that are resident on the CPU. The programmer can access memory on a device without explicit copies. If a memory access is part of a page that is not on the device, the UVM-driven triggers a page-fault that automatically migrates the page to the requesting device. The UVM driver can also evict pages from a given device back to host memory when the total paged memory size exceeds device memory (NVIDIA, 2023a).
UVM provides several benefits in regard to programmability. First, programmers are exposed to a single unified address space that they can access as if the whole allocated chunk of memory is resident on a single GPU. Any copies occurring around the system are implicit and hidden from the programmer’s view. Additionally, UVM allows memory oversubscription whereby more memory can be allocated than all the multi-GPU device memory combined. This is possible since most of the memory can stay on the CPU and be paged in whenever a given device requests it (NVIDIA, 2021; Shao et al., 2022).
@@ 3.2. GPUDirect Technologies
Introduced in CUDA 3.1, GPUDirect 1.0 allowed GPUs and NICs to share the same pinned memory region. Prior, the pinned memory regions in system memory for GPUs and the NIC were separate. By implication, to communicate GPU data across nodes, the GPU first copied the data to its pinned memory region, the CPU then copied it to the NIC’s memory region, only then it can be accessed by the NIC, which sent it across the network, as shown in Figure 2 1. The intermediate CPU-initiated copy from GPU $\rightarrow$ NIC pinned memory regions adds CPU overhead and increases the latency for GPU communication. GPUDirect 1.0 introduced a shared memory GPU-NIC pinned memory region, thus, avoiding the intermediate CPU-initiated copy (Rossetti et al., 2016; Shainer et al., 2011).
@@
Along with the introduction of UVA, the CUDA 4.0 release added support for direct peer-to-peer communication among GPUs in a single node that share the same PCIe root complex (NVIDIA, 2011). This functionality was encapsulated in a technology known as GPUDirect 2.0 or GPUDirect P2P. Instead of staging data through the host, GPUs could now directly access each other’s memory over PCIe, establishing, for the first time, a direct GPU-to-GPU data path. These changes led to two new communication mechanisms: P2P DMA Copies whereby a cudaMemcpy call would trigger a DMA transfer directly between source and target GPU memories and P2P Direct Load / Stores using which the GPUs could directly access data by dereferencing pointers to the remote GPU buffers. GPUDirect P2P also added support for NVLink (Section 3.4) when the latter technology was introduced (NVIDIA, 2023g, 2012; Rossetti et al., 2016).
GPUDirect P2P provided two main benefits. It eliminated redundant GPU $\leftrightarrow$ CPU copies and host buffers, which were required when the transfers were staged through the CPU. Also, by eliding the need to maintain communication buffers on the host and providing a new communication mechanism (P2P Direct Load / Stores), GPUDirect P2P increased the convenience of multi-GPU programming (NVIDIA, 2012).
We note that P2P DMA Copies can also work without UVA support. If UVA is not enabled, P2P DMA Copies can be performed using the cudaMemcpyPeer() variants by explicitly specifying the target GPU. However, P2P Direct Load / Stores will not work without UVA as directly accessing a remote GPU’s pointer presumes a unified address space (NVIDIA, 2023a).
@@ 3.2.3. GPUDirect RDMA
With the introduction of GPUDirect RDMA in CUDA 5.0, direct communication between NVIDIA GPUs across nodes became feasible. GPUDirect RDMA facilitates a direct communication channel between GPUs and third-party devices through standard PCIe features. The technology exposes segments of GPU memory on the PCIe memory resource, referred to as the Base Address Register (BAR) region. This enables NICs to directly read/write GPU memory without routing through the host (NVIDIA, 2023d). Analogously, AMD offers ROCm RDMA (previously called ROCnRDMA) (AMD, 2023d, b).
GPUDirect RDMA provides several optimizations to the data path, namely by eliminating additional copies to host memory, reducing inherent latencies stemming from GPU-NIC interaction, increasing bandwidth and reducing CPU overhead. Support for GPUDirect RDMA has been integrated into several leading communication libraries including GPU-aware MPI implementations, NCCL and NVSHMEM.

### v3

@@ 1. Introduction
In recent years, GPUs have become the accelerator of choice for a vast array of applications across both HPC and ML. This quick adoption, driven mainly by the GPU’s massive parallelism and high memory bandwidth, means that most of modern cloud and HPC capability is now concentrated in clusters of GPUs. As of November 2025, 9 of the 10 leading Top500 supercomputers rely on GPU clusters for acceleration and this trend is likely to continue (top500). The only system, Fugaku, without GPUs employs a highly vectorized CPU architecture combined with High-Bandwidth Memory.
While using scores of GPUs has been shown to significantly accelerate computation, communication between the GPUs can quickly become a scalability bottleneck (Peta-scale-stencil) (padal2). Traditionally, multi-GPU communication, both within and across nodes, had always been the responsibility of the CPU. From their conception, GPUs were thought of as devices that can supply large amounts of computation but that are also inherently dependent on the CPU for auxiliary tasks like communication. In this CPU-centric model of execution, the routines relaying data for consumption by the GPUs are oblivious to the GPUs’ existence.
@@
  • In Section 2, we define the terminology, provide a definition for GPU-centric communication. We taxonomize and abstract the existing approaches within a node and across nodes to eliminate any confusion.
  • In Section 3, we present the history and discuss the vendor-provided mechanisms for enabling communication and networking, managing memory across devices in a multi-GPU execution. These mechanisms are used as building blocks for higher-level GPU-centric software libraries.
  • In Section 4, we list and compare the main communication libraries for both intra- and inter-node setups, discuss their benefits and challenges and rely on existing benchmarking works to provide insights into their performance.
@@
Some methods and technologies discussed in this survey are inherently tied to proprietary ecosystems. While a substantial portion of the available literature and deployed systems focuses on NVIDIA technologies, we present a balanced perspective by incorporating corresponding mechanisms from other vendors, such as AMD and Intel, whenever information is available. At the same time, several capabilities remain exclusive to specific vendors; we explicitly identify these vendor-specific features to ensure clarity and completeness.
@@ 2. Terminology and Communication Types
We can loosely define GPU-centric communication as mechanisms that reduce the involvement of the CPU in the critical path of multi-GPU execution. This is a very broad definition covering a wide spectrum of solutions, involving both the vendor-level improvements that grant GPUs autonomy in communication and user-level implementations that leverage those improvements. To make this distinction clear, we discuss them in separate sections. In Section 3, we focus on communication mechanisms and primitives provided natively as part of the NVIDIA CUDA and AMD ROCm runtimes. In Section 4, we discuss how these mechanisms give rise to higher-level, GPU-centric communication libraries and present both vendor-provided software support by AMD, Intel and NVIDIA as well as other industry and academic solutions.
@@ 2.1. Intra-Node Communication
Type API Data Path Examples
3 Device Native Device Device NVSHMEM / ROCSHMEM / Intel SHMEM, GPU-side, direct load/store (P2P)
4 Host Fallback Device Host Direct load/store (No P2P)
Table 1. Types of intra-node communication methods
The communication method in 1, referred as host native, is made on the host side and does not involve direct P2P (peer-to-peer) access between the devices. These involve all methods that can be launched on the host side with P2P access disabled. Otherwise, as 2 (host-controlled) shows, the communication does not involve an extra copy to the host memory, and passes directly through PCIe, NVLink or Infinity Fabric interconnects. GPUCCL (i.e., NCCL, RCCL, oneCCL, etc), GPU-aware MPI and *memcpy operations have host-side API, so they may belong to both 1 and 2. By enabling direct access to peer device memory, the device-side API eliminates CPU involvement from both the data and control paths in intra-node communication, as illustrated by 3, referred to as device native, in Figure 1. NVSHMEM, ROCSHMEM, Intel SHMEM offer host-side API as well, but require P2P access, so their host-side APIs belong to 2 and device-side API is of type 3. In-kernel P2P direct load and stores offer similar functionality and belong to type 3, but work even with P2P access disabled, which is type 4, where data path falls back to the host.
2.2. Inter-Node Communication
@@ 1 Host Native Host Host Host Through host (2 copies) GPU-Aware MPI before 2
2 Pinned Host Native Host Host Host Through host (1 copy) GPUDirect 1.0 in GPU-Aware MPI, NCCL, RCCL, oneCCL
3 GPU RDMA D/H Host Host Direct GPUDirect RDMA, ROCmRDMA in GPU-Aware MPI, NCCL, NVSHMEM, ROCSHMEM
4 GPU-Triggered D/H Host Device Depends on 3 GPUDirect Async in GPU-Aware MPI, NCCL v2.28, NVSHMEM, ROCSHMEM
@@ 5 Device Native Device Device Device Depends on 3 NVSHMEM with IBGDA, GPUrdma
We identify five main categories when classifying inter-node communication. From 1 to 5, as Table 2 and Figure 2 show, more components of communication calls were moved to the GPU side, while the data transfer path saw a reduction in the number of data copies required to reach the NIC. Over the years, the communication methods and the corresponding technologies depict the optimizations of both data transfer and communication control, which will be discussed in Section 3. First, entirely CPU-side communication method in 1 was available, which has been improved by removal of extra copy between CPU-GPU and CPU-NIC buffers in 2 with the help of shared pinned memory between GPU and NIC. This reduced the latency for GPU-NIC transfers. After that, GPU RDMA (Remote Direct Memory Access) 3 facilitated direct access to the GPU memory from NIC over PCIe minimizing the data path between them. 4 represents GPU-triggered communication technologies such as GPUDirect Async and GPU-TN (GPU-TN-GPU-Triggered-Networking-for-Intra-Kernel-Communication), where the GPU has become capable to initiate communications, given that CPU prepared the packets on the NIC in advance. 5 moves the packet preparation and interaction with the NIC entirely to the GPU as well, making device-native communication possible.
The types given in Table 2 and Figure 2 do not reflect all possible combinations, since some libraries, based on the configuration and the available hardware may lead to different combinations of data paths and control. For example, without the RDMA technology, even with GPU-side communication control, the data path will involve the host memory. For instance, Intel SHMEM uses a proxy thread on the host to execute actual RDMA using a standard OpenSHMEM library even though the communication call is issued within a device kernel running on the GPU.
@@ 3. Vendor Mechanisms
In this section we discuss the vendor-provided mechanisms for enabling communication and networking, managing memory across devices in a multi-GPU execution. These mechanisms are provided by GPU programming model runtimes or as part of the extended APIs. The technologies are classified into four categories: memory managers, GPUDirect technologies, hardware, and libraries. Figure 3 summarizes the technologies provided by NVIDIA, detailing their timeline and availability.
Next, we introduce the memory management mechanisms and GPUDirect technologies, followed by the hardware support that served as precursors and ultimately made these communication methods viable. These technologies form the backbone of the higher-level GPU-centric libraries, which are discussed in Section 4.
@@ 3.1.1. Page-Locked / Pinned Memory
By default, memory allocated on the host using device malloc (e.g. cudaMalloc(), hipMalloc() is pageable and is not accessible to the GPU. When a transfer between pageable host memory and device memory is performed, the GPU runtime must first stage the host data through a temporary buffer in page-locked memory and then copy the data from page-locked memory to the GPU. To avoid the pageable $\rightarrow$ page-locked memory copy, cudaMallocHost() allows allocating page-locked memory directly, skipping the intermediate copy stage. Because of this, page-locked memory is also referred to as zero-copy or pinned memory (CUDA-Programming-Guide-Release-12.2; CUDA-Blog-Pinned-Memory-How-to-Optimize-Data-Transfers-in-CUDA-C/C++).
Pinned memory, known for its high bandwidth and low latencies in host-device transfers (Evaluating-Characteristics-of-CUDA-Communication-Primitives-on-High-Bandwidth-Interconnects), efficiently coordinates CPU-GPU execution by enabling direct system-wide access. It is also utilized with GPUDirect RDMA for improved inter-node communication (Evaluating-Modern-GPU-Interconnects). However, its physical memory locking can lead to high memory consumption, potentially impacting system performance with excessive allocations (CUDA-Blog-Pinned-Memory-How-to-Optimize-Data-Transfers-in-CUDA-C/C++).
@@ 3.1.2. Unified Virtual Addressing (UVA)
UVA is a memory management technique which allows all GPUs and CPUs within a node to share the same unified virtual address space (CUDA-Programming-Guide-Release-12.2; CUDA-4.0-Release-Notes). Prior to UVA, host $\leftrightarrow$ device and device $\leftrightarrow$ device copies had to explicitly specify the direction of transfer. With UVA, the physical memory location can be inferred from pointer values, thus, reducing the overhead of managing separate memory spaces and enabling libraries to simplify their interfaces (Talk-Peer-to-Peer-and-Unified-Virtual-Addressing).
3.1.3. IPC
In early GPU runtimes versions, pointers could not be accessed across process boundaries, so memory copies between GPU buffers had to go through the host, creating a bottleneck. To overcome this limitation, Inter-Process Communication (IPC) enables processes on the same node to access device buffers of other processes without additional copies (CUDA-4.1-Release-Notes). With IPC, memory handles are created and passed between processes using standard IPC mechanisms, resulting in lower latencies than staging copies through the host.
@@ 3.1.4. Unified Virtual Memory (UVM)
UVM allows for the allocation of managed memory through cudaMallocManaged() calls by creating a single address space accessible to all processors within a single node. UVM works by dividing the requested memory into pages that are resident on the CPU. The programmer can access memory on a device without explicit copies. If a memory access is part of a page that is not on the device, the UVM-driven triggers a page-fault that automatically migrates the page to the requesting device. The UVM driver can also evict pages from a given device back to host memory when the total paged memory size exceeds device memory (CUDA-Programming-Guide-Release-12.2).
UVM provides several benefits in regard to programmability. First, programmers are exposed to a single unified address space that they can access as if the whole allocated chunk of memory is resident on a single GPU. Any copies occurring around the system are implicit and hidden from the programmer’s view. Additionally, UVM allows memory oversubscription whereby more memory can be allocated than all the multi-GPU device memory combined. This is possible since most of the memory can stay on the CPU and be paged in whenever a given device requests it (CUDA-Blog-Improving-GPU-Memory-Oversubscription-Performance; Oversubscribing-GPU-Unified-Virtual-Memory-Implications-and-Suggestions).
@@ 3.2. GPUDirect Technologies
GPUDirect 1.0 allowed GPUs and NICs to share the same pinned memory region. Prior, the pinned memory regions in system memory for GPUs and the NIC were separate. By implication, to communicate GPU data across nodes, the GPU first copied the data to its pinned memory region, the CPU then copied it to the NIC’s memory region, only then it can be accessed by the NIC, which sent it across the network, as shown in Figure 2 1. The intermediate CPU-initiated copy from GPU $\rightarrow$ NIC pinned memory regions adds CPU overhead and increases the latency for GPU communication. GPUDirect 1.0 introduced a shared memory GPU-NIC pinned memory region, thus, avoiding the intermediate CPU-initiated copy (Talk-State-Of-GPUDirect-Technologies; GPUDirect-1.0-The-development-of-Mellanox/NVIDIA-GPUDirect-over-Infiniband).
@@
Along with the introduction of UVA, the CUDA 4.0 release added support for direct peer-to-peer communication among GPUs in a single node that share the same PCIe root complex (CUDA-4.0-Release-Notes). This functionality was encapsulated in a technology known as GPUDirect 2.0 or GPUDirect P2P. Instead of staging data through the host, GPUs could now directly access each other’s memory over PCIe, establishing, for the first time, a direct GPU-to-GPU data path. These changes led to two new communication mechanisms: P2P DMA Copies whereby a cudaMemcpy call would trigger a DMA transfer directly between source and target GPU memories and P2P Direct Load / Stores using which the GPUs could directly access data by dereferencing pointers to the remote GPU buffers. GPUDirect P2P also added support for NVLink (Section 3.4) when the latter technology was introduced (NVIDIA-GPUDirect; Talk-NVIDIA-GPUDirect; Talk-State-Of-GPUDirect-Technologies).
GPUDirect P2P provided two main benefits. It eliminated redundant GPU $\leftrightarrow$ CPU copies and host buffers, which were required when the transfers were staged through the CPU. Also, by eliding the need to maintain communication buffers on the host and providing a new communication mechanism (P2P Direct Load / Stores), GPUDirect P2P increased the convenience of multi-GPU programming (Talk-NVIDIA-GPUDirect).
We note that P2P DMA Copies can also work without UVA support. If UVA is not enabled, P2P DMA Copies can be performed using the cudaMemcpyPeer() variants by explicitly specifying the target GPU. However, P2P Direct Load / Stores will not work without UVA as directly accessing a remote GPU’s pointer presumes a unified address space (CUDA-Programming-Guide-Release-12.2).
@@ 3.2.3. GPUDirect RDMA
With the introduction of GPUDirect RDMA in CUDA 5.0, direct communication between NVIDIA GPUs across nodes became feasible. GPUDirect RDMA facilitates a direct communication channel between GPUs and third-party devices through standard PCIe features. The technology exposes segments of GPU memory on the PCIe memory resource, referred to as the Base Address Register (BAR) region. This enables NICs to directly read/write GPU memory without routing through the host (GPUDirect-RDMA). Analogously, AMD offers ROCm RDMA (previously called ROCnRDMA) (ROCnRDMA; ROCK-Kernel-Driver). GPUDirect RDMA provides several optimizations to the data path, namely by eliminating additional copies to host memory, reducing inherent latencies stemming from GPU-NIC interaction, increasing bandwidth and reducing CPU overhead.